# Understanding LLM Embeddings for Regression

### v1

@@ 1 Introduction and Related Work
One understudied case however is the use of service-based LLM embeddings - fixed vector representations derived from pre-trained (but frozen) language models, which are ubiquitously offered among most LLM services (OpenAI, 2023; Google, 2024). Although they are used frequently in recent applications such as retrieval (Karpukhin et al., 2020), semantic similarity (Li et al., 2020), and a variety of other downstream language tasks (Liu et al., 2020), there has been very little research for their general and direct use in regression, outside of specific applications such as Bayesian Optimization (Nguyen et al., 2024).
In contrast to decoding-based regression techniques, embedding-based regression allows the possibility of cheap data-driven training using inexpensive and customizable post-embedding layers such as multi-layer perceptrons (MLPs). However, as shown in Figure 1, when the domain of a simple function is expressed using high-dimensional embeddings, unexpected characteristics and irregularities can arise, prompting the need for a thorough analysis. Furthermore, LLMs by default are not explicitly trained for embedding-based regression, rather purely for token generation, and thus it is worth analyzing the emergent behaviors of LLMs when applied to embedding-based regression.
This paper investigates the behavior of these LLM embeddings when used as features for traditional regression tasks. Most notably, our findings are:
@@ 2.2 Regression Tasks
For obtaining data, we may either sample $(x,y)$ pairs (in the case of synthetic objectives where $x$ are uniformly sampled from $\mathcal{X}$), or use the given offline data (in the case of real-world tasks, where they were actual evaluations from an optimization trajectory), using a standard 8-1-1 train-validation-test split.
Due to the inherent differing of metric scales across tasks, it would be inappropriate to aggregate results based on scale-dependent metrics such as mean squared error (MSE). Furthermore, we found that the selection of the regression metric (e.g. Kendall-Tau, Pearson, mean squared error, mean absolute error) did not matter for comparisons, as they all strongly correlated with each other. Thus, by default we report the Kendall-Tau ranking correlation, which is always within $[0,1]$ and can also be aggregated across different tasks.
@@ 3.1 High Dimensional Regression
We begin by demonstrating cases in which LLM embeddings better represent inputs over high degree-of-freedom spaces than traditional representations. In Figure 2, we show that for a subset of functions, LLM embeddings possess surprising robustness, retaining the same performance for varying DOFs whereas traditional baselines such as XGBoost and MLPs significantly falter over higher DOFs.
Figure 2: Higher ($\uparrow$) is better. Degrees of freedom (DOF) vs Kendall-Tau correlation for various BBOB functions. Results are averaged over 12 runs for each regression method. Each task’s data consists of 500 $(x,y)$ evaluations sampled uniformly across the search space, using a 8-1-1 split for train-validation-test.
This result is not universal however, as we show in Appendix A.1, this pattern does not apply for a few selected functions, but nonetheless it occurs in the majority of the BBOB functions. We further corroborate this observation over real-world tasks in Table 1. We see that in general, regressions on LLM embeddings outperform traditional methods more often for tasks with higher DOFs (AutoML and XLA).
@@ 3.2 LLM Embedding Smoothness
Figure 3: Left-skewness ($\leftarrow$) is better. NLFDs induced by $\phi_{\text{LLM}}$ (Gemini Nano) and $\phi_{\text{trad}}$. Top: Cases where $\phi_{\text{LLM}}$ outperforms $\phi_{\text{trad}}$ for regression. Bottom: Vice-versa where $\phi_{\text{trad}}$ outperforms $\phi_{\text{LLM}}$.
@@
    Normalize each embedding vector $\phi(x)$ coordinate-wise to have zero mean and unit variance across the dataset $\mathcal{D}$.
@@ Appendix C Example String Representations
Below are example string representations of $x$ for different regression task families. ‘…’ denotes that there are actually more parameters, but we omit them due to page limits.
@@
AutoML batch_size:128, ml_feature_selection_threshold:0.05, model_type:‘DNN_ESTIMATOR’, activation_fn:‘selu’, batch_norm:‘False’, bucketization_strategy:‘mdl’, dropout:0.071, hidden_units:359
Init2Winit lr_hparams.base_lr:0.0696, opt_hparams.0.hps.one_minus_b1:0.2823, opt_hparams.0.hps.one_minus_b2:0.0432, opt_hparams.1.hps.weight_decay:0.0023
XLA auto_cross_replica_sharding:‘False’, rematerialization_percent_shared_memory_limit:97, spmd_threshold_for_windowed_einsum_mib:100000, ...

### v2

@@ 1 Introduction and Related Work
One understudied case however is the use of service-based LLM embeddings - fixed vector representations derived from pre-trained (but frozen) language models, which are ubiquitously offered among most LLM services (OpenAI, 2023; Google, 2024; Anthropic, 2024). Although they are used frequently in recent applications such as retrieval (Karpukhin et al., 2020), semantic similarity (Li et al., 2020), and a variety of other downstream language tasks (Liu et al., 2020), there has been very little fundamental research around their use in regression, outside of specific applications such as Bayesian Optimization (Nguyen et al., 2024; Kristiadi et al., 2024).
In contrast to decoding-based regression techniques, embedding-based regression allows the possibility of cheap data-driven training using inexpensive and customizable post-embedding layers such as multi-layer perceptrons (MLPs). However, as shown in Figure 1, when the domain of a simple function is expressed using high-dimensional embeddings, unexpected characteristics and irregularities can arise, prompting the need for a thorough analysis. Furthermore, LLMs by default are not explicitly trained for embedding-based regression, rather purely for token generation, and thus it is worth analyzing the emergent behaviors of LLM embeddings when applied to regression.
This paper investigates the behavior of these LLM embeddings when used as features for standard tabular regression tasks. Most notably, our findings are:
@@ 2.2 Regression Tasks
For obtaining data, we may either sample $(x,y)$ pairs (in the case of synthetic objectives where $x$ are sampled from $\mathcal{X}$), or use the given offline data (in the case of real-world tasks, where they were actual evaluations from an optimization trajectory), using a standard 8-1-1 train-validation-test split.
Due to the inherent differing of metric scales across tasks, it would be inappropriate to aggregate results based on scale-dependent metrics such as mean squared error (MSE). Furthermore, we found that the selection of the regression metric (e.g. Kendall-Tau, Pearson, mean squared error, mean absolute error) did not matter for comparisons, as they all strongly correlated with each other. Thus, by default we report the Kendall-Tau ranking correlation, which is always within $[0,1]$ and can be aggregated across different tasks.
@@ 3.1 High Dimensional Regression
We begin by demonstrating cases in which LLM embeddings better represent inputs over high DOF spaces than traditional representations. In Figure 2, we show that for a subset of functions, LLM embeddings possess surprising robustness, retaining the same performance for varying DOFs whereas traditional baselines such as XGBoost and MLPs significantly falter over higher DOFs.
Figure 2: Higher ($\uparrow$) is better. Degrees of freedom (DOF) vs Kendall-Tau correlation for various BBOB functions. Results are averaged over 12 runs for each regression method. Each task’s data consists of 500 $(x,y)$ evaluations sampled uniformly across the input space, using a 8-1-1 split for train-validation-test.
This result is not universal however, as we show in Appendix A.1, this pattern does not apply for some selected functions, but nonetheless it occurs in the majority of the BBOB functions. We further corroborate this observation over real-world tasks in Table 1. We see that in general, regressions on LLM embeddings outperform traditional methods more often for tasks with higher DOFs (AutoML and XLA).
@@ 3.2 LLM Embedding Smoothness
Figure 3: Left-skewness ($\leftarrow$) is better. NLFDs induced by $\phi_{\text{LLM}}$ (T5-XXL) and $\phi_{\text{trad}}$. Top: Cases where $\phi_{\text{LLM}}$ outperforms $\phi_{\text{trad}}$ for regression. Bottom: Vice-versa where $\phi_{\text{trad}}$ outperforms $\phi_{\text{LLM}}$.
@@
    Full-batch normalize, i.e. apply shifting and scaling to each $\phi(x)$ so that in aggregate, $\mathcal{D}$ has zero mean and unit variance per coordinate.
@@ A.2 Real World Results
Despite Table 1 of the main body showing that there were numerous cases where LLM embeddings outperform traditional ones, we remind the reader in Table 11 that on average, LLM embeddings still slightly underperform.
@@ Appendix C Example String Representations
Table 4 contains example string representations of $x$ for different regression task families.
@@
AutoML batch_size:128, ml_feature_selection_threshold:0.05, model_type:’DNN_ESTIMATOR’, activation_fn:’selu’, batch_norm:’False’,
bucketization_strategy:’mdl’,dropout:0.071, hidden_units:359
Init2Winit lr_hparams.base_lr:0.0696, opt_hparams.0.hps.one_minus_b1:0.2823,
opt_hparams.0.hps.one_minus_b2:0.0432,
opt_hparams.1.hps.weight_decay:0.0023
XLA auto_cross_replica_sharding:’False’,
rematerialization_percent_shared_memory_limit:97,
spmd_threshold_for_windowed_einsum_mib:100000, ...
@@
Table 4: Example $x$ representations from each of the regression task families. ‘…’ denotes that there are actually more parameters, but we omit them due to length.